# Basis-Spline Assisted Coded Computing: Strategies and Error Bounds

### v1

@@ III-D Function Reconstruction using Spline
where $\{B_{j,p}(z)\}$ are basis functions of B-spline, and the coefficients $\{C_{j,p}\}$ are obtained using an appropriate interpolation method using the $N$ function evaluations $\{f(u(\alpha_{i}))\mid i=0,1,\ldots,N-1\}$. Using (5), the desired function $f(\mathbf{X}_{j})$ are reconstructed by evaluating the spline at the points $\{\beta_{j}\}_{j=0}^{K-1}$ as $\mathbf{Y}_{j}=r_{B-spline}(\beta_{j})$. Note that these are approximate evaluations, i..e, $\mathbf{Y}_{j}\approx f(\mathbf{X}_{j})$, for $j\in[K-1]$, and the accuracy of this reconstruction depends on the reconstruction accuracy of $r_{B-spline}(z)$ in (5). In the next section, we present our interpolation method to choose the set $\{C_{j,p}\}$.
III-E Clamped B-Spline Interpolation for BSCC
In our framework, we employ cubic B-splines, i.e., $p=3$, to approximate $f(u(z))$. Let $\{B_{j,3}(z)\}_{j=0}^{N+1}$ denote the set of cubic B-spline basis functions, where the spline space $\mathcal{S}_{3,t}\triangleq\operatorname{span}\{B_{0,3}(z),B_{1,3}(z),\ldots,B_{N+1,3}(z)\},$ is constructed over the knot vector $t$ given by
$t=\bigl[\underbrace{\alpha_{0},\ldots,\alpha_{0}}_{4},\alpha_{1},\alpha_{2},\ldots,\alpha_{N-2},\underbrace{\alpha_{N-1},\ldots,\alpha_{N-1}}_{4}\bigr],$ (6)
where $\{\alpha_{i}\}_{i=0}^{N-1}$ are the avaulation points for BSCC. Note that the above knot vector corresponds to a clamped (open) knot vector for cubic B-splines. In general, for splines of degree $p$, the extreme knot points are repeated $p$ times. As a result, in our case, $\alpha_{0}$ and $\alpha_{N-1}$ are replicated thrice. Given $N$ function evaluations $\{f(u(\alpha_{i}))\mid i=0,1,\ldots,N-1\}$ obtained from the worker nodes at a set of distinct and ordered evaluation points $\{\alpha_{i}\}_{i=0}^{N-1}$, we approximate the composite function $f(u(z))$ by a cubic spline $s(z)\in\mathcal{S}_{3,t}$ of the form
@@ III-E Clamped B-Spline Interpolation for BSCC
Consequently, the resulting linear system is underdetermined and admits infinitely many solutions. To obtain a unique solution, p1=2p-1=2 additional constraints are required. Therefore, we impose the natural spline boundary conditions $s^{\prime\prime}(\alpha_{0})=0\quad\text{and}\quad s^{\prime\prime}(\alpha_{N-1})=0,$ where $s^{\prime\prime}(z)$ denotes the double derivative of $s(z)$. Since $s^{\prime\prime}(z)=\sum_{j=0}^{N+1}C_{j,3}\,B^{\prime\prime}_{j,3}(z),$ these boundary conditions yield two additional linear constraints on the spline coefficients [14]. Evaluating at the boundary points gives
@@
By augmenting the rectangular system with these two boundary constraints, we obtain a square B-spline matrix $\mathbf{B}\in\mathbb{R}^{(N+2)\times(N+2)}$. The resulting linear system can be written as $\mathbf{B}\mathbf{c}=\bar{\mathbf{f}}$, where $\bar{\mathbf{f}}=[0~\mathbf{f}^{T}~0]^{T}$ and
$\mathbf{B}=\begin{bmatrix}B^{\prime\prime}_{0,3}(\alpha_{0})&B^{\prime\prime}_{1,3}(\alpha_{0})&\cdots&B^{\prime\prime}_{N+1,3}(\alpha_{0})\\ B_{0,3}(\alpha_{0})&B_{1,3}(\alpha_{0})&\cdots&B_{N+1,3}(\alpha_{0})\\ B_{0,3}(\alpha_{1})&B_{1,3}(\alpha_{1})&\cdots&B_{N+1,3}(\alpha_{1})\\ \vdots&\vdots&\ddots&\vdots\\ B_{0,3}(\alpha_{N-1})&B_{1,3}(\alpha_{N-1})&\cdots&B_{N+1,3}(\alpha_{N-1})\\ B^{\prime\prime}_{0,3}(\alpha_{N-1})&B^{\prime\prime}_{1,3}(\alpha_{N-1})&\cdots&B^{\prime\prime}_{N+1,3}(\alpha_{N-1})\end{bmatrix}$. (10)
The two additional zero entries in the right-hand side vector correspond to the imposed natural spline boundary conditions. Before solving this linear system, we next present results on the structure of $\mathbf{B}$ and establish results on its invertibility as well as the computational complexity of its inversion.
@@
For cubic B-spline i.e., $p=3$, the B-spline matrix $\mathbf{B}\in\mathbb{R}^{(N+2)\times(N+2)}$ in (10) is banded and full rank, and hence the associated linear system $\mathbf{B}\mathbf{c}=\bar{\mathbf{f}}$ admits a unique solution.
@@
For cubic B-splines, the matrix $\mathbf{B}\in\mathbb{R}^{(N+2)\times(N+2)}$ in (10) has upper and lower bandwidth $2$. Subsequently, the associated linear system can be solved using banded LU decomposition with computational complexity $\mathcal{O}(8(N+2))$, which is smaller than that of a dense matrix.
Overall, after obtaining the spline coefficients $\{C_{j,3}\}_{j=0}^{N+1}$, the desired function values are reconstructed as explained in Section III.
IV Experimental Results
@@
Figure 1: Average relative error (in dB) of the BSCC and BACC decoding schemes with Lagrange and Berrut encoding for $f(\mathbf{X})=\mathbf{X}\sin(\mathbf{X})$ and $f(\mathbf{X})=\frac{1}{1+e^{-\mathbf{X}}}$, with parameters $N=100$, $K=8$, and $\mathbf{X}\in\mathbb{R}^{40\times 5}$.
The plots in Fig. 1 confirm that BSCC achieves a remarkable improvement in approximation accuracy over BACC for non-polynomial functions, particularly when the number of stragglers is small, and exhibits a significant improvement even when the number of stragglers increases. Moreover, the results highlight that, to maintain a relative error on the order of $10^{-2}$ or $10^{-3}$, the BSCC framework can tolerate a substantially larger number of stragglers compared to the BACC framework.

### v2

@@ III-D Function Reconstruction using B-spline
where $\{B_{j,p}(z)\}$ are B-spline basis functions of degree $p$, and the coefficients $\{C_{j,p}\}$ are obtained using an appropriate interpolation method using the $N$ function evaluations $\{f(u(\alpha_{i}))\mid i=0,1,\ldots,N-1\}$. Using (3), the desired function $f(\mathbf{X}_{j})$ are reconstructed by evaluating the spline at the points $\{\beta_{j}\}_{j=0}^{K-1}$ as $\mathbf{Y}_{j}=r_{B-spline}(\beta_{j})$. Note that these are approximate evaluations, and the accuracy of this reconstruction depends on the reconstruction accuracy of $r_{B-spline}(z)$ in (3). In the next section, we present our interpolation method to choose the set $\{C_{j,p}\}$ using cubic B-splines, i.e., p=3p=3.222In approximation theory, it is well known that among various degrees of B-spline, cubic B-splines achieve a good balance between the accuracy of approximation and computational complexity for interpolation.
III-E Cubic B-Spline Interpolation for BSCC
With the use of cubic B-splines, i.e., p=3\!p=\!3, let $\{B_{j,3}(z)\}_{j=0}^{N+1}$ denote the set of cubic B-spline basis functions, where the spline space $\mathcal{S}_{3,\mathbf{t}}$ is constructed over the $4$-clamped knot vector $\mathbf{t}$ given by
$\mathbf{t}=\bigl[\underbrace{\alpha_{0},\ldots,\alpha_{0}}_{4},\alpha_{1},\alpha_{2},\ldots,\alpha_{N-2},\underbrace{\alpha_{N-1},\ldots,\alpha_{N-1}}_{4}\bigr],$ (4)
where $\{\alpha_{i}\}_{i=0}^{N-1}$ are the distinct and ordered evaluation points for BSCC. Given $N$ function evaluations $\{f(u(\alpha_{i}))\mid i=0,1,\ldots,N-1\}$, we approximate the composite function $f(u(z))$ by a cubic spline $s(z)\in\mathcal{S}_{3,\mathbf{t}}$ of the form
@@ III-E Cubic B-Spline Interpolation for BSCC
Consequently, the resulting linear system is underdetermined and admits infinitely many solutions. To obtain a unique solution, p1=2p\!-1\!=2 additional constraints are required. Therefore, we impose the natural spline boundary conditions $s^{\prime\prime}(\alpha_{0})=0\quad\text{and}\quad s^{\prime\prime}(\alpha_{N-1})=0,$ where $s^{\prime\prime}(z)$ denotes the double derivative of $s(z)$. Since $s^{\prime\prime}(z)=\sum_{j=0}^{N+1}C_{j,3}\,B^{\prime\prime}_{j,3}(z),$ these boundary conditions yield two additional linear constraints on the spline coefficients [15]. Evaluating them gives
@@
By augmenting the rectangular system with these two boundary constraints, we obtain an $(N+2)\times(N+2)$ square B-spline matrix. The resulting system of linear equations can be written as $\mathbf{B}\mathbf{c}=\bar{\mathbf{f}}$, where $\bar{\mathbf{f}}=[0~\mathbf{f}^{T}~0]^{T}$ and
$\mathbf{B}=\begin{bmatrix}\mathbf{a}\\ \mathbf{B}_{\mathrm{rect}}\\ \mathbf{b}\end{bmatrix}$, (8)
where the row vectors $\mathbf{a},\mathbf{b}\in\mathbb{R}^{1\times(N+2)}$are given by
$\mathbf{a}=\bigl[B^{\prime\prime}_{0,3}(\alpha_{0}),\;B^{\prime\prime}_{1,3}(\alpha_{0}),\;\ldots,\;B^{\prime\prime}_{N+1,3}(\alpha_{0})\bigr],$
$\mathbf{b}=\bigl[B^{\prime\prime}_{0,3}(\alpha_{N-1}),\;B^{\prime\prime}_{1,3}(\alpha_{N-1}),\;\ldots,\;B^{\prime\prime}_{N+1,3}(\alpha_{N-1})\bigr].$
The two additional zero entries in $\bar{\mathbf{f}}$ correspond to the imposed natural spline boundary conditions. Before solving this linear system, we next present results on the structure of $\mathbf{B}$ and establish results on its invertibility as well as the computational complexity for its inversion.
@@
[…]
Proof.
See Appendix -A for the proof. ∎
@@
For cubic B-splines, the matrix $\mathbf{B}\in\mathbb{R}^{(N+2)\times(N+2)}$ in (8) has upper and lower bandwidth of $2$. Subsequently, the associated linear system can be solved using banded LU decomposition with computational complexity $\mathcal{O}(8(N+2))$[17], which is smaller than that of a dense matrix.
Overall, after obtaining the spline coefficients $\{C_{j,3}\}_{j=0}^{N+1}$, the desired function values are reconstructed as explained in Section III-D. In the rest of the section, we discuss the modifications required for interpolation in the presence of stragglers, i.e., when $S>0$. Let $\mathcal{F}\!\!=\!\!\{l_{1},l_{2},\ldots,l_{M}\}$ denote the index set of the M=NSM\!\!=\!\!N-S non-stragglers. Let $\{\mathbf{R}_{l_{1}},\mathbf{R}_{l_{2}},\ldots,\mathbf{R}_{l_{M}}\}$ denote the corresponding results received at the master, where $\mathbf{R}_{l_{i}}=f\!\left(u(\alpha_{l_{i}})\right)$ for $l_{i}\in\mathcal{F}$. Let $\{\alpha_{l_{i}}\}_{i=0}^{M-1}\subset\{\alpha_{i}\}_{i=0}^{N-1}$ denote the ordered set of evaluation points corresponding to the non-stragglers in the set $\mathcal{F}$. To generate the B-spline basis functions, the master constructs the associated $4$-clamped knot vector from evaluation points of the non-stragglers set. Subsequently, $\mathbf{B}_{rect}$ defined in (6) reduces to the dimension $M\times(M+2)$, where the basis functions $\{B_{i,3}(z)\}_{i=0}^{M+1}$ are evaluated at the points $\{\alpha_{l_{i}}\}_{i=0}^{M-1}$. Further, after imposing the boundary conditions on $\mathbf{B}_{rect}$, we obtain $\mathbf{B}$ defined in (8) of dimension $(M+2)\times(M+2)$. Finally, solving the resulting linear system yields the spline coefficients, which are subsequently used to recover the desired values $\{f(\mathbf{X}_{j})\}_{j=0}^{K-1}$. Similar to BACC, there is no strict recovery threshold for BSCC as well. The master reconstructs the function using the evaluations from the non-stragglers, with accuracy improving as more evaluations are received.
IV Error Bounds for BSCC
As discussed in Section III-D, the reconstruction accuracy of $\{f(\mathbf{X}_{j})\}_{j=0}^{K-1}$ depends on the accuracy with which $r_{B-spline}(z)$ approximates $f(u(z))$. Therefore, we are interested in deriving bounds on the approximation error $r_{B-spline}(z)-f(u(z))||_{\infty}$, where $||\cdot||_{\infty}$ denotes the infinity norm over the support set. Henceforth, we use $g(z)$ to refer to $f(u(z))$ and $Tg(z)$ to refer to $r_{B-spline}(z)$, assuming that $T$ is the underlying B-spline interpolation operator that approximates $g(z)$ to $Tg(z)\in\mathcal{S}_{3,\mathbf{t}}$. Towards evaluating the interpolator operator $T$, we recollect this definition from [11].
Definition 5.
Let $C[a,b]$ represent the set of all continuous functions defined over $[a,b]$. Further, for any function $g(z)\in C[a,b]$, let $Tg(z)\in\mathcal{S}_{3,\mathbf{t}}$ denote the spline function obtained by applying the operator $T$ to $g(z)$. Then the operator norm of $T$ is defined as
$\|T\|:=\sup_{g(z)\neq 0}\frac{\|Tg(z)\|_{\infty}}{\|g(z)\|_{\infty}}.$ (9)
For the proposed BSCC interpolation with cubic B-splines, we derive an upper bound on the operator norm $T||$, which is presented in the following proposition.
Proposition 3.
For a given $N$, consider the BSCC setting when $S=0$. Let $T$ denote our spline interpolation operator that maps the composite function $g(z)$ to a spline function $Tg(z)\in\mathcal{S}_{3,\mathbf{t}}$. Then, $\|T\|$ is upper bounded as
$\|T\|\leq\|[\mathbf{B}^{-1}]_{\mathrm{sub}}\|_{\infty},$ (10)
where $\mathbf{B}^{-1}]_{\mathrm{sub}}\!\in\!\mathbb{R}^{(N+2)\times{N}}$ denotes the submatrix of $\mathbf{B}^{-1}$ after excluding the first and the last columns, where $\mathbf{B}$ is given in (8).
Proof.
See Appendix -B for the proof. ∎
For a given set of distinct ordered evaluation points $\{\alpha_{i}\}_{i=0}^{N-1}$, the above upper bound $\|[\mathbf{B}^{-1}]_{\mathrm{sub}}\|_{\infty}$ can be computed using (8) in a straightforward manner. However, in the following theorem, we derive an upper bound on $\|[\mathbf{B}^{-1}]_{\mathrm{sub}}\|_{\infty}$ only as a function of the knot spacings.
Theorem 1.
For a given $N$, a knot vector $\mathbf{t}$, when S=0S\!\!\!=\!\!\!0, let $h_{\max}\!\!\!=\!\!\!\max_{\begin{subarray}{c}0\leq i\leq N-2\\ \end{subarray}}\left(t_{i+1}-t_{i}\right)$ and $h_{\min}\!\!=\!\!\min_{\begin{subarray}{c}0\leq i\leq N-2\end{subarray}}\left(t_{i+1}-t_{i}\right)$ represent the maximum and minimum knot spacing between consecutive distinct knots in the knot vector. Then
$\|[\mathbf{B}^{-1}]_{\mathrm{sub}}\|_{\infty}\leq C_{1}(N+2)\frac{h_{\max}}{h_{\min}},$
where $C_{1}>0$ is a constant, which depends only on the degree of the spline.
Proof.
See Appendix -C for the proof. ∎
Further, using the above upper bound on $\|[\mathbf{B}^{-1}]_{\mathrm{sub}}\|_{\infty}$, we use the results from [11] to present an upper bound on the approximation error of BSCC.
Theorem 2.
For a given $N$, consider the BSCC scheme when $S=0$. Then, for every function $g(z)\in C[a,b]$, the interpolation error of BSCC satisfies
$\|g(z)-Tg(z)\|_{\infty}$ $\leq\Bigl(1+C_{1}(N+2)\frac{h_{\max}}{h_{\min}}\Bigr)D_{S_{3,t}}(g(z)),$ (11)
where $D_{S_{3,t}}(g(z))=\inf_{s(z)\in S_{3,\mathbf{t}}}\|g(z)-s(z)\|_{\infty}$, and all other constants are as defined in Theorem 1.
Proof.
See Appendix -D for the proof. ∎
While the above theorem gives a bound on the approximation error of BSCC for an arbitrary function $g(z)$, there is no simpler representation of $D_{S_{3,t}}(z)$ in general. However, for a special class of functions, i.e., when $g(z)\in C^{4}[a,b]$, where $C^{4}[a,b]$ denotes the set of all functions over $[a,b]$ whose fourth derivative is also continuous, we use the results in [11, Chapter XII, page 149] to bound $D_{S_{3,t}}(g(z))$, and subsequently derive an upper bound on (11).
Corollary 1.
For the BSCC setting considered in Theorem 2, if $g(z)\in C^{4}[a,b]$, then there exists a constant $C>0$, depending only on the degree of spline, such that
$\|g(z)-Tg(z)\|_{\infty}\leq C\bigg(1+C_{1}(N+2)\frac{h_{\max}}{h_{\min}}\bigg)\,h_{max}^{4}\|g^{\prime\prime\prime\prime}(z)\|_{\infty},$ (12)
where $g^{\prime\prime\prime\prime}(z)$ represents the fourth derivative of $g(z)$ in $[a,b]$, and all other constants are as defined in Theorem 2.
All the above results are presented when $S=0$. However, with $S>0$, Theorem 2 and Corollary 1 continue to hold with the exception that $N$ must be replaced by NSN-S and the parameters $h_{max}$ and $h_{min}$ must be computed using the evaluation indices of non-stragglers.
IV-A Error Bounds Comparison with BACC
Although BSCC allows any choice of evaluation points $\{\alpha_{i}\}_{i=0}^{N-1}$ over $[a,b]$, one can select $\{\alpha_{i}\}_{i=0}^{N\!-\!1}$ as Chebyshev nodes of the second kind, defined by $\alpha_{i}=\cos\!\left(\frac{i\pi}{N}\right)$, for $i\in[N]$. In such a case, with $S$ stragglers, the upper bound in (12) reduces to
$2C\left(\sin^{4}\!\left(\frac{(S+1)\pi}{2N}\right)+\frac{C_{1}(N-S+2)}{h_{min}}\sin^{5}\!\left(\frac{(S+1)\pi}{2N}\right)\right)\|g^{\prime\prime\prime\prime}(z)\|_{\infty},$ (13)
where $h_{min}$ is independent of $S$, and is known to follow $\mathcal{O}(\frac{1}{N^{2}})$ behavior, for the class of Chebyshev nodes. We recall from [9, Theorem 9] that the approximation error of BACC varies with $S$ and $N$ as
$\bigg(1+\frac{(1+S)(3+S)\pi^{2}}{4}\bigg)\sin\!\left(\frac{(S+1)\pi}{2N}\right),$ (14)
Figure 1: Average relative error (in dB) of the BSCC and BACC schemes with Lagrange and Berrut encoding with N=100N\!=\!100, $K=8$.
after excluding all other constants. Provided $g(z)\in C^{4}[a,b]$, using (13) and (14), we make the following inferences: With $S=0$, the error bound of BACC varies as $\mathcal{O}(\frac{1}{N})$, whereas the error bound of BSCC varies as $\mathcal{O}(\frac{1}{N^{2}})$, thereby indicating faster decay of approximation error with the number of workers. For $0<S<<N$, in the regime of large $N$, the error bound of BACC behaves as $\mathcal{O}(\frac{S^{3}}{N})$, whereas the error bound of BSCC varies as $\mathcal{O}(\frac{S^{5}}{N^{2}})$. This indicates that, for small $S$ and large $N$, the absolute error values of BSCC are expected to be much lower than that of BACC owing to the inverse-quadratic behavior with respect to $N$. Furthermore, along the lines of BACC, for a fixed $N$, we observed that the error bound of BSCC also grows with respect to $S$. In the next section, we attempt to validate these observations using experimental results for arbitrary functions.
V Experimental Results
@@
VI Future work
In this work, we have proposed a B-spline-assisted coded computing framework applicable to distributed computation of arbitrary functions. Our future work is to extend the proposed framework to incorporate Byzantine workers, develop robust encoding strategies based on B-splines, and adopt the proposed method to accelerate distributed computational tasks in a wide range of applications, including distributed learning.